find the value of x woth the help of below equation 
x^2 + 2 = 6 .

Grade:6

4 Answers

CHEKURI NIKHIL
7 years ago
dear sumit  the given equation X^2+2=6 can be solved as follows:
x^2=6-2
x^2=4
x^2=2^2
x=2. so the vavue of x is 2.
i didnot understand the woth in your question
thanks
g v subba reddy
7 years ago
given equation is x2+2=6
adding -2 on both sides
we will get x2=4
now the value of x becomes +2 or -2
This is so because by squaring +2 or -2 we will get the same result that is 4
So for this reason we will take values of x as +2 or-2
Siddharth Acharya
7 years ago
X2+2=6
=> X2=4
=> square root(x2) = square root(4)
=> |x|=4
because whenever we remove root sign, we add a modulus. Thus we have,
=> x= +2 or -2
imran
7 years ago
Hi SUMIT
THE VALUE OF X IS
WE CAN WRITE THE EQUATION X^2+2=6
THAT IMPLISE TO X^2=6-2
THAT EQUALS TO X^2=4
THEN X=SQUARE ROOT OF 4
THEREFORE THE VALUE OF X IS TO BE +2 AND -2
                                           OK THANK Q